OpenStudy (anonymous):

Find w(4) for the function w(x) = 3x + 7.

OpenStudy (anonymous):

w(4)=3*4+7=19

OpenStudy (anonymous):

3*4+7=19

OpenStudy (amistre64):

yep, thats all there is to it; multiply, then add ....

OpenStudy (anonymous):

just replace the variable of x with the value of 4, which is given in parantasises :D w(4) = 3*4 + 7